# LLM as Effective Streaming Processor: Bridging Streaming-Batch Mismatches with Group Position Encoding

### v1

@@ Abstract
Junlong Tong1,2, Jinlan Fu3, Zixuan Lin4, Yingqi Fan2, Anhao Zhao2, Hui Su5, Xiaoyu Shen2thanks: Corresponding author, 1Shanghai Jiao Tong University 2Ningbo Key Laboratory of Spatial Intelligence and Digital Derivative, Institute of Digital Twin, EIT 3National University of Singapore 4University of Science and Technology of China 5Meituan Inc. jl-tong@sjtu.edu.cn     xyshen@eitech.edu.cn
@@ 1 Introduction
Figure 1: Two streaming paradigms of LLMs: (a) Batch-streaming simulates batch-processing, while interleaved-streaming encodes streaming data in arrival order. (a-1) Input-Attention Mismatch: Whether the source can attend to the target. (a-2) Out-Attention Mismatch: Whether the target can attend to the latest source. (a-3) Position-ID Mismatch: Whether the position IDs reflect the true token order. (b) Batch-streaming relies on re-encoding to mimic batch-processing, differing only in accessible information range.
@@
  • Out-Attention Mismatch: Batch-streaming allows decoded output tokens to attending to all received input tokens by KV cache re-encoding, while interleaved-streaming mode limits each output token’s attention to the subset of inputs available at decoding time.
@@
Building on the identification of these mismatches, we systematically studied their effects on LLM performance. Our analysis revealed that input-attention mismatch does affect streaming model performance. In contrast, out-attention and position-ID mismatches have negligible effects. A common assumption is that streaming models require re-encoding of previously generated content to mitigate token position inconsistencies arising from the incremental nature of streaming setting Raffel et al. (2024); Guo et al. (2024a); He et al. (2024), as shown in Figure 1(b). However, our empirical findings do not support this hypothesis. Instead, we observe that re-encoding the output is not necessary  [^1] We clarify that re-encoding the target tokens is solely for refining the generation of the latest token without altering previously generated content.. This discrepancy with the common assumption raises a fundamental question: How does position encoding impact LLMs in streaming scenarios? And how should we design appropriate position encoding for streaming LLMs?
@@ 2 Streaming-Batch Mismatches
Out-Attention Mismatch
In batch-streaming mode, each generated output token $y_{k}$ can attend to all input tokens $\mathbf{X}$ by KV cache re-encoding:
@@ 3 Impact Analysis of Mismatches
Table 1: The BLEU performance variations reflect the stepwise elimination of mismatches between batch processing and streaming. Interleaved-streaming represents the presence of all three mismatches. Batch-streaming (No re.) corresponds to batch-streaming with interleaved position encoding, where the input-attention mismatch is eliminated. Batch-streaming (Pos re.) further removes the position-ID mismatch through position re-encoding. Finally, Batch-streaming (All re.) eliminates the out-attention mismatch by re-encoding the KV cache.
@@
The interleaved-streaming mode, which exhibits all three mismatches, serves as our baseline for comparison. The batch-streaming mode eliminates input-attention mismatch by preventing source tokens from attending to generated target tokens. Building on this, we apply the same positional encoding as interleaved-streaming within the batch-streaming framework. Notably, without KV cache and position embedding re-encoding, the batch-streaming approach still retains both out-attention mismatch and position-ID mismatch.
@@
Re-encoding can address the remaining two mismatches. We further decompose re-encoding into two components: KV cache re-encoding and position embedding re-encoding. The former enables target tokens to attend to the most recently available tokens, thereby resolving the out-attention mismatch. The latter corrects the position-ID mismatch by adjusting position embeddings to align with the streaming paradigm. Expanding on this, the batch-streaming paradigm with position embedding re-encoding further resolves the position-ID mismatch while still retaining out-attention mismatch.
@@
Effects of Out-Attention Mismatch
The KV cache re-encoding can address the remaining out-attention mismatch. On the setting of the former, incorporating both KV cache and position embedding re-encoding into batch-streaming paradigm eliminates all mismatches, making it closely resemble the batch-processing setting.
@@ 4 Impact Analysis of Position Encoding
Table 3: Performance comparison of different models with various wait-k policies and target start IDs. $\Delta$ represents the range of variation in BLEU scores when the start id of target token takes different values. We use bold to indicate the smallest variation. Underline represents the largest variation.
@@ 4.1 Is Position Encoding Necessary for Streaming Tasks?
4.2 Group Position Encoding Is An Option for Streaming Tasks.
@@ Appendix F Model Efficiency
The results in the figure show that the proposed grouped-streaming paradigm eliminates the need for re-encoding, resulting in significant throughput improvements: over 5 $\times$ speedup under the wait-9 setting and more than 11 $\times$ speedup under wait-5, compared to the re-encoding baseline.
@@ G.2 Example of Streaming Decoding
Figure 8 is an example of streaming decoding process.
Figure 8: English to German streaming translation. The bold indicate the most recently read or generated content.

### v2

@@ Abstract
Junlong Tong1,2,3, Jinlan Fu4, Zixuan Lin5, Yingqi Fan3, Anhao Zhao3, Hui Su6, Xiaoyu Shen2,3thanks: Corresponding author 1Shanghai Jiao Tong University 2Ningbo Key Laboratory of Spatial Intelligence and Digital Derivative 3Institute of Digital Twin, EIT  4National University of Singapore 5University of Science and Technology of China 6Meituan Inc. jl-tong@sjtu.edu.cn     xyshen@eitech.edu.cn
@@ 1 Introduction
Figure 1: Two streaming paradigms of LLMs: (a) Batch-streaming simulates batch-processing, while interleaved-streaming encodes streaming data in arrival order. (a-1) Input-Attention Mismatch: Whether the source tokens can attend to the target tokens. (a-2) Output-Attention Mismatch: Whether the target tokens can attend to the new source token. (a-3) Position-ID Mismatch: Whether the position IDs reflect the actual token order. (b) Batch-streaming relies on (b-1) KV cache re-encoding and (b-2) position re-encoding to simulate batch-processing.
@@
  • Output-Attention Mismatch:Batch-streaming allows decoded output tokens to attending to all received input tokens by KV cache re-encoding, while interleaved-streaming mode limits each output token’s attention to the subset of inputs available at decoding time.
@@
Building on the identification of these mismatches, we systematically studied their effects on LLM performance. Our analysis revealed that input-attention mismatch does affect streaming model performance. In contrast, output-attention and position-ID mismatches have negligible effects. A common assumption is that streaming models require re-encoding of previously generated content to mitigate token position inconsistencies arising from the incremental nature of streaming setting Raffel et al. (2024); Guo et al. (2024a); He et al. (2024), as shown in Figure 1(b). However, our empirical findings do not support this hypothesis. Instead, we observe that re-encoding the output is not necessary  [^1] We clarify that re-encoding the target tokens is solely for refining the generation of the latest token without altering previously generated content.. This discrepancy with the common assumption raises a fundamental question: How does position encoding impact LLMs in streaming scenarios? And how should we design appropriate position encoding for streaming LLMs?
@@ 2 Streaming-Batch Mismatches
Output-Attention Mismatch
In the batch-streaming mode, each generated output token $y_{k}$ can attend to all input tokens $\mathbf{X}$ by KV cache re-encoding:
@@ 3 Impact Analysis of Mismatches
Table 1: The BLEU performance variations reflect the stepwise elimination of mismatches between batch processing and streaming. Interleaved-streaming represents the presence of all three mismatches. Batch-streaming (No re.) corresponds to batch-streaming with interleaved position encoding, where the input-attention mismatch is eliminated. Batch-streaming (Pos re.) further removes the position-ID mismatch through position re-encoding. Finally, Batch-streaming (All re.) eliminates the output-attention mismatch by re-encoding the KV cache.
@@
The interleaved-streaming mode, which exhibits all three mismatches, serves as our baseline for comparison. The batch-streaming mode eliminates input-attention mismatch by preventing source tokens from attending to generated target tokens. Building on this, we apply the same positional encoding as interleaved-streaming within the batch-streaming framework. Notably, without KV cache and position embedding re-encoding, the batch-streaming approach still retains both output-attention mismatch and position-ID mismatch.
@@
Re-encoding can address the remaining two mismatches. We further decompose re-encoding into two components: KV cache re-encoding and position embedding re-encoding. The former enables target tokens to attend to the most recently available tokens, thereby resolving the output-attention mismatch. The latter corrects the position-ID mismatch by adjusting position embeddings to align with the streaming paradigm. Expanding on this, the batch-streaming paradigm with position embedding re-encoding further resolves the position-ID mismatch while still retaining output-attention mismatch.
@@
Effects of Output-Attention Mismatch
The KV cache re-encoding can address the remaining output-attention mismatch. On the setting of the former, incorporating both KV cache and position embedding re-encoding into batch-streaming paradigm eliminates all mismatches, making it closely resemble the batch-processing setting.
@@ 4 Impact Analysis of Position Encoding
Table 3: Performance comparison of different models with various wait-k policies and target start IDs. $\Delta$ represents the range of variation in BLEU scores when the start id of target token takes different values. We use bold to indicate the smallest variation and underline to represent the largest variation.
@@ 4.1 Is Position Encoding Necessary for Streaming Tasks?
4.2 Group Position Encoding Is An Option for Streaming Tasks
@@ 7 Conclusion
Acknowledgements
We thank EIT and IDT High Performance Computing Center for providing computational resources for this project. This work was supported by the 2035 Key Research and Development Program of Ningbo City under Grant No.2024Z127.
@@ Appendix F Model Efficiency
The results in the figure show that the proposed grouped-streaming paradigm eliminates the need for re-encoding, resulting in significant throughput improvements: over 5 $\times$ speedup under the wait-9 setting and more than 11 $\times$ speedup under wait-5 setting, compared to the re-encoding baseline.
@@ G.2 Example of Streaming Decoding
Figure 8 is an example of streaming reading and decoding process.
Figure 8: An example on wait-5 reading/writing policy. The bold indicate the most recently content.